OpenStudy (anonymous):

hellp Choose the equation of the line passing through the point (-2, 4) and perpendicular to y = -one halfx + 4.

OpenStudy (anonymous):

y = -x/2 + 4 Slope = -1/2 Since the lines are perpendicular the Product of their slopes = -1 -1/2 * m1 = -1 m1 = 2 Nw using point slope Form (-2, 4) y - 4 = 2 (x+2) y = 2x+8

OpenStudy (anonymous):

thaNK YOU wanna help with mor yahoo or any one

OpenStudy (anonymous):

help Wake Up coffee shop offers two different sizes of coffee. The small coffee costs $2 and the grande coffee costs $3. It is your turn to buy coffee for the office and the bill came to $18. This scenario can be modeled by the equation 2s + 3g = 18. Which is not a possible combination of small and grande coffees?

OpenStudy (aripotta):

are there choices?

OpenStudy (anonymous):

yes hold on one min

OpenStudy (anonymous):

9 small and 1 grande 0 small and 6 grande 3 small and 4 grande 6 small and 2 grande

OpenStudy (anonymous):

i think it is c

OpenStudy (aripotta):

it's not c

OpenStudy (anonymous):

fail lol

OpenStudy (aripotta):

it gives you this: 2s + 3g = 18. so for each choice, just substitute and see if it's true or not. the choice that's not true is your answer

OpenStudy (aripotta):

for c, 2(3) + 3(4) = 18 6 + 12 = 18 18 = 18, which is true. so it couldn't be c

OpenStudy (anonymous):

oh ok never mind then

OpenStudy (anonymous):

i came out with a

OpenStudy (aripotta):

yup

OpenStudy (anonymous):

can u help me with some more i am really bad at math and i am doing practice problems

OpenStudy (aripotta):

i'll probably help with one more, but i'm really tired lol. it's 7:25am...haven't slept :l

OpenStudy (anonymous):

ohh my you better go sleep i have another person that can help me and just listend to ur thing

OpenStudy (anonymous):

its so funny lol

OpenStudy (aripotta):

:D

OpenStudy (anonymous):

hope you have a good sleep lol byye

OpenStudy (aripotta):

night! or morning! or whatever! bai and good luck :)

OpenStudy (anonymous):

lol byyye anD @Shadowys

OpenStudy (anonymous):

LOL here! :)

OpenStudy (anonymous):

Choose the equation of the line passing through the point (4, -2) and perpendicular to y = one thirdx - 7. y = -3x - 14 y = 3x - 10 y = -3x + 10 y = 3x + 14

OpenStudy (anonymous):

hahahah thank you for helping me my mother is mad i have a 63 im math lol eeeecc

OpenStudy (anonymous):

lol that's pretty good actually. since it is perpendicular, then \(m_1=\frac{1}{3}\) \(m_2=-3\) invert and add neg. passing through (4, -2) \(y+2 = -3(x-4)\) which gives \(y=-3x +10\) you follow?

OpenStudy (anonymous):

yes ok then

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

oo yes ok then

OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 3x - 3y = 4. slope = -1; y intercept at (0, -four thirds) slope = 1; y intercept at (0, -four thirds) slope = 3; y intercept at (0, -1) slope = -3; y intercept at (0, 1)

OpenStudy (anonymous):

slope being gradient, so we rearrange that eq. -3y=-3x+4 \(y=x-\frac{4}{3}\) so, slope is 1, y intercept is (\(\frac{-4}{3}\),0)

OpenStudy (anonymous):

oh i did that backward lol hahah ok i have 3 more lol

OpenStudy (anonymous):

ok Identify the correct x intercept and y intercept of the equation x - 2y = 8. x intercept = (8, 0); y intercept = (0, 4) x intercept = (8, 0); y intercept = (0, -4) x intercept = (-8, 0); y intercept = (0, -4) x intercept = (-2, 0); y intercept = (0, 8)

OpenStudy (anonymous):

i think it is b or a

OpenStudy (anonymous):

well, when they ask for y intercept, it means the value when x=0. So, is it b or a? try plugging in x=0 for the y intercept

OpenStudy (anonymous):

now im coming out with 2 0 and o 8 negitive

OpenStudy (anonymous):

hmm? when x=0, the eq becomes -2y=8 then y=-4

OpenStudy (anonymous):

ohhhhhhh smap never mind oh wow ok lol i skiped like 2 things lol

OpenStudy (anonymous):

for this one i came out with a but im not sure its right Choose the equation of the line passing through the point (1, -5) and parallel to y = 2x + 4. y = -2x + 7 y = 2x - 3 y = 2x - 7 y = -2x + 3

OpenStudy (anonymous):

lol i think you got it?

OpenStudy (anonymous):

yayy and one more idk what it is tho

OpenStudy (anonymous):

Choose the equation of the line passing through the point (-4, -2) and parallel to y = 1 fourthx + 2. y = 4x - 1 y = 4x - 3 y = 1 fourthx - 3 y = 1 fourthx - 1

OpenStudy (anonymous):

\(y=\frac{1}{4} x +2\) so the gradient of the next line will also be \(\frac{1}{4}\) what did you get after applying the formula?

OpenStudy (anonymous):

y 4x-

OpenStudy (anonymous):

opps 1

OpenStudy (anonymous):

eh? lol

OpenStudy (anonymous):

am i right lol

OpenStudy (anonymous):

or is it 3

OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of -6 and a y-intercept of 9.

OpenStudy (anonymous):

heres the last one

OpenStudy (anonymous):

Hmm that one isn't perpendicular so actually it should be 4. \(y+2=\frac{1}{4} (x+4)\) \(y=\frac{1}{4}x+1-2)\) \(y=\frac{1}{4}x-1\)

OpenStudy (anonymous):

ohh ok and did u see the last question

OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of -6 and a y-intercept of 9.

OpenStudy (anonymous):

this would be : y=-6x +9 you follow?

OpenStudy (anonymous):

ohhhh lol ok

OpenStudy (anonymous):

THANK YOU THANK YOU YOYOUOYOYOUOYOUOYOUO LOL THANK YOU

OpenStudy (anonymous):

lol you're welcome :) glad to.

OpenStudy (anonymous):

hahahah

OpenStudy (anonymous):

well umm do you have time to help with more or are u buissy

OpenStudy (anonymous):

lol actually I have.

OpenStudy (anonymous):

you have to go

OpenStudy (anonymous):

or u have time

OpenStudy (anonymous):

time~ lol lol dun worry. fire away

OpenStudy (anonymous):

ok fisrt one

OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (-2, -1) and has a slope of 5. y - 1 = 5(x - 2) y + 1 = 5(x + 2) y + 2 = 5(x + 1) y - 2 = 5(x - 1)

OpenStudy (anonymous):

ya im virtual schooled

OpenStudy (anonymous):

i think its c

OpenStudy (anonymous):

hmm, you've got it reversed though, it's the other way around. It's b.

OpenStudy (anonymous):

ohh ok next one lol hold on

OpenStudy (anonymous):

Choose the equation of the horizontal line that passes through the point (4, -3). y = 4 x = -3 x = 4 y = -3

OpenStudy (anonymous):

s a

OpenStudy (anonymous):

opps a

OpenStudy (anonymous):

well, the x must equal 4, or y equal -3. once again, you've got it reversed? lol

OpenStudy (anonymous):

the coords is (x,y)=(4,-3)

OpenStudy (anonymous):

so its b

OpenStudy (anonymous):

and it's horizontal, so it should be a y=something. it's y=-3. note the coords

OpenStudy (anonymous):

this one again i think is a Which of the following is the slope between the two points (0, -4) and (-2, 5)? -nine halves two ninths -two ninths nine halves

OpenStudy (anonymous):

yup! this is a

OpenStudy (anonymous):

ya yyy lol another one iss lol i think it is a or c again hahah Choose the slope-intercept equation of the line that passes through the point (-2, 2) and is parallel to y = 4x + 7.

OpenStudy (anonymous):

lol options?

OpenStudy (anonymous):

heyyyy

OpenStudy (anonymous):

y = 4x + 10 y = one fourthx + 5 halves y = -4x - 6 y = -one fourthx + 3 halves

OpenStudy (anonymous):

sorry my computer stoped working

OpenStudy (anonymous):

lol yh mine too actually well, considering that it is parallel, then the gradient must be 4 too. the only option with 4 as gradient is...?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

opps a

OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, -6) and is perpendicular to y = 3x - 6. y = one thirdx - 8 y = -3x + 12 y = 3x - 24 y = -one thirdx - 4

OpenStudy (anonymous):

i think it is d

OpenStudy (anonymous):

yup, it's a., and yup, it's d. LOL i think you've got it!

OpenStudy (anonymous):

ok another one is ,,,,,,,,

OpenStudy (anonymous):

now this one i think is c but i really dont knwo Choose the correct procedure used to graph the equation 3x + 9y = -18 using the intercepts method. Solve the equation for y to get a y-intercept of (0, -2) and a slope of -one third. Plot the y-intercept, count the slope to get another point and draw the line that passes through the two points. Choose two different values; plug them into the equation to find the corresponding y-values. Plot the two points and draw the line that passes through the two points. Let x equal zero and solve for y to find a y-intercept of (0, -2). Let y equal zero and solve for x to find an x-intercept of (-6, 0). Plot the two points and draw a line passing through the two points. Solve the equation for y to find an intercept of (-2, 0); solve the equation for x to find a y-intercept of (0, -6). Plot the two points and draw a line passing through the two points.

OpenStudy (anonymous):

i really dont get this one lol

OpenStudy (anonymous):

Well, the intercept form actually means number three lol

OpenStudy (anonymous):

ohhhhhhhhh ok nvm next one i got it Identify the correct slope and y-intercept of the equation 5x - 3y = -6 slope = -5; y-intercept = (0, -6) slope = five thirds; y-intercept = (0, 2) slope = -three fifths; y-intercept = (0, 2) slope = five thirds; y-intercept = (0, -6)

OpenStudy (anonymous):

i think its d

OpenStudy (anonymous):

you've got the slope right, but the y-intercept can't be -6

OpenStudy (anonymous):

ohhhhhhhhhhh c then right

OpenStudy (anonymous):

ok next one Choose the equation of the vertical line that passes through the point (-5, 9). y = -5 y = 9 x = -5 x = 9